OpenStudy (osiris):

find theapproximate length of the hypotenuse of a right triangle with leg lenghts 8.4 cm and 7.6 cm

OpenStudy (lgbasallote):

hypotenuse = \(\sqrt{a^2 + b^2}\) hypotenuse = \(\sqrt{(8.4)^2 + (7.6)^2}\) use calculator to solve

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8.4^2 + 7.6^2 = h^2 }\)